Find the value of x if a^(x-2)[(a^(2x+2) + a^(1-x)] = a^(-3)[(a^9 + a^2)]
Answer by stanbon(75887)

Find the value of x if Find the value of x if a^(x-2)[(a^(2x+2) + a^(1-x)] = a^(-3)[(a^9 + a^2)]
-----
Left side simplified:::a^(x-2)[(a^(2x+2) + a^(1-x)] = a^[x-2+2x+2] + a^[x-2+1-x] = a^3x+a^-1
and
Right side simplified:: a^(-3)[(a^9 + a^2)] = a^6 + a^-1
----
So, 3x = 6 and x = 2
-----
Cheers,
Stan H.
-----------